# Distributed Multisensor ISAC

### v1

@@ IV-A Multipath Propagation Channel and Multipath Exploitation
Multipath propagation issues are crucial to the performance of mobile radio and are critical to system design. Multipath is responsible for channel dispersion in delay, Doppler, and angular domain. It leads to temporal variability, fast and slow fading. It has been considered the most limiting feature of system performance that needs to be mitigated by equalization, precoding and link adaptation in different domains. Despite a certain diversity gain advantage, multipath originally was seen as harmful for high data rates and reliable communication. The situation has dramatically changed with the advent of MIMO radio access, which has shown that multipath can improve communication and enhance capacity in complicated, multipath-rich environments. Also in radio localization, multipath may turn from a foe to a friend if correctly exploited [ref55]. Therefore, measurement and model based propagation studies in rich scattering environments are of outstanding importance also for ISAC system design [ref3]. Although the topic would deserve more attention, we can give only a concise overview on the most relevant propagation effects to support understanding of basic MS ISAC design issues.
@@
The consideration of the propagation channel for radar localization differs remarkably from the situation in communication. Generally speaking, in radar the desired information is contained in the geometric structure of the propagation, while in communication the information is encoded in the unknown transmitted symbols, which have to be received correctly. With ISAC, where communication and sensing functions are integrated, we even have a mixture of both which casts different views on the influence of propagation. Figure 6 depicts the geometric structure of a single bistatic multipath propagation channel. In communication, the total energy per symbol sent and received is used for symbol recognition, hence for maximizing mutual information. In ISAC only those paths that are routed via the target can contribute to target detection and localization. However, the LOS path between Tx and Rx (which is often misleadingly referred to simply as “interference”) is also important. In addition to its contribution to communication, it helps for Tx2Rx frame synchronization, submits a reference for differential eToF and eDoppler estimation in bistatic setups and correlation reference signal recovery, which is necessary for CPCL target estimation as described above. In contrast, all other propagation paths that are not routed via the target are useless for radar sensing, perhaps disturbing and therefore referred to as clutter. The role of the target channel, which consists of all multiple paths that are routed via the target, can be thought of as a relay channel as highlighted in Figure 6.
@@
The target is most clearly identified by the two LOS paths (Tx-to-target and target-to-Rx in Figure 6), which we call the Bistatic Target Path (BTP). However, the simple triangular LOS-related model would be ruined if one or more of the three LOS paths are blocked. While the blocked direct Tx2Rx LOS must be given special consideration since it delivers the correlation reference, there are several possibilities to exploit target related multipath. In general, the associated multipath links that are also routed via the target additionally contribute to targets illumination and visibility and may reveal relevant geometric information for localization even if the respective LOS is missing. Spatial precoding and dedicated estimation measures would be necessary to exploit this multipath interaction (“multipath exploitation”), may it be enhanced visibility through spatial multipath focusing (increased illumination power) on the target or enabling target visibility even when the target LOS to Rx is obstructed. Here we can only give a very concise overview highlighting effects and indicating respective measures to mitigate or exploit multipath.
@@
Proactive exploitation of multipath goes further ahead. While precoding and associated resource allocation measures are shortly referred to in the next section, a detailed discussion about multipath exploiting procedures is beyond of this paper. Roughly speaking, there seem to be two classes of multipath exploiting estimation methods. One is called “finger printing”. It is based on calculating some position related features from observed ISAC data. In the most simple case, this can be received power which undergoes spatial selective fading. However, differential delays, Doppler or directions that can be better attributed to relative target position and movement can be also used. While explicit geometrical calculation of position from these parameters may not be possible in multipath rich environment, machine learning can be applied to identify targets position from measured fingerprints. In addition, reference information calculated from a ray tracing simulation of the scenario (a “digital twin”) can serve as a learning data base [ref56, ref57].
Another approach consists in explicitly embedding a multipath propagation data model based on ray tracing into the position estimation loop. Assuming we know the positions of the dominant interaction points of the environment, the associated reflections can be used as further (or auxiliary) spatially orthogonal measurements like additional illuminations generated by virtual anchor nodes. This inverse model-based estimator effectively traces back the observed specular data to the unknown target position [ref51]. The advantage relative to finger printing is less learning effort which is required (the explicit knowledge about the structure of the environment is directly applied) and we have more control about which interacting objects are included. This may be also important if “artificial interactions” generated by Reconfigurable Intelligent Surfaces (RISs) or by amplify-and-forward relays are used.
Yet another approach that enables target localization with obstructed LOS from Tx to or from the target to Rx (e.g. if the target gets lost when tracking) can be used for indirect localization if we have knowledge about the effect of obstruction and especially if we explicitly know the position of the obstructing object [ref58, ref59]. Even when the target itself is obscuring the Tx2Rx LOS we can detected the presence of the target when it moves through the LOS between Tx and Rx (forward scattering radar [ref60]).
@@
where $H_{\rm los}$, $H_{\rm dti}$, $H_{\rm dtr}$, $H_{\rm sti}$, $H_{\rm str}$, and $H_{\rm clt}$ denote the FRFs for the LOS, direct target illumination, direct target response, scattered target illumination, scattered target response, and clutter, respectively. Clutter are all contribution that are not related to the target (except LOS). Whether the scattered target illumination and response should be considered a “clutter” depends on whether it is exploited or not, i.e., if it constitutes unwanted interference.
@@ IV-B Multistatic Target Reflectivity and Diversity Gain
The target’s contribution to the end-to-end multipath routing from transmitter to receiver is described by its Multistatic Reflectivity Function (MRF), which is the bistatic transfer function of all pairs of incoming and outgoing BTP waves including target-related LOS and multipath. Therefore, the target deserves special attention in ISAC propagation measurement and modeling. A complete MRF description would be based on 2x2 full polarimetric bistatic angular scanning. The MRF measurement range described in [ref3] uses wideband illumination in order to characterize target’s inherent time variability (e.g. because of rotating wheels or propellers in case of cars or drones or moving arms and legs in case of walking persons), which will cause characteristic micro Doppler patterns. Together with the bistatic spread in delay and angle, we obtain a full MRF description that enables advanced propagation studies, recognition of extended targets and high performance ray optical modeling in multipath rich environments. Note that these characteristics are distance dependent as they are influenced by spherical wavefronts.
@@
At greater distances and lower bandwidths (which does not allow to resolve the shape of the target), the advantage of distributed MIMO radar shifts to spatial diversity, which affects the probability of target detection. The use of spatial diversity for mitigation of fading is already well known from MIMO communication where it is related to the signal at the receiver antenna position. In distributed MIMO radar it is related the target. Since radar targets are usually larger than several wavelengths of the radio carrier (“electrically large targets”) and complex in terms of many structural details of small size, even cavities, the scattered field is composed of contributions from a great number of seemingly isolated scattering centers, with the resulting superposition being a complex circular symmetric Gaussian variable (iid in real and imaginary part). Therefore, the reflected power of the target (related to its Radar Cross Section (RCS)) undergoes a chi-square distribution with two degrees of freedom. This applies as long as the multiple reflections make a similar contribution (in terms of power) to the resulting target backscatter. In case we have a dominating (specular) contribution, the distribution will change to a non-central chi-square distribution (a squared Rician distribution). Moreover, the reflected power maybe subjected to a slow variability over time. These scintillations are responsible for signal fading, which has influence to target detection statistics. In radar technology, these statistical variations are combined in so called Swerling’s target models [ref32]. In mobile radio communications, the same problem is known as small scale and large scale fading.
@@
The diversity gain increases the probability of detection in case of multiple observations [ref10c]. It depends on the decorrelation of the combined observations and can be characterized as cooperative diversity or macro-diversity. Therefore, distributed MIMO radar is often referred to as statistical MIMO radar to distinguish it from MIMO radar with coherent signal combining. The possible diversity gain depends of course on the number of independent observations, which in the best case can be up to $N^{2}$ (see section above). A pleasant side effect is that the claimed distributed spatial structure is also advantageous for estimating the 3D dynamic state vector which will be described in the next section. The influence of the aspect angle is also evident when we consider special target shapes, such as those targets that attempt to conceal themselves by minimizing their monostatic RCS, also known as stealth targets. At the same time, we have to emphasize a very specific spatial constellation of Tx and Rx, the so-called forward scattering case where Tx and Rx face each other from opposite directions and the target is in between. Surprisingly enough, this situation has many advantages for target detection as the received power may undergo a clear increase because of the creeping wave along the targets surface and the resulting lens effect. For higher frequencies (mmWave) and depending on the distance between target and antennas, also shadowing may occur. The latter case may become relevant for short distance sensing in car-to-car scenarios.
@@
While the combined chi-square distributions apply to multiple observations of electrically large targets, we need to distinguish the case of extended targets. These are targets that are resolved in the delay domain and are therefore represented by a small number of neighboring range bins. For a given target, this situation depends on the bandwidth of the ISAC waveform. The more bandwidth we have, the more range bins we can resolve –- for the same target. Since the observed reflectivity response of the neighboring range bins is not independent because it originates from the same target seen from the same direction, we can try to maximize received power, respectively the detection probability, by matching the transmit signal to the target response (target matched filter).
@@ V MS ISAC Access and Estimation
In Section III, we describe a very general architecture of MS MIMO ISAC systems. According to the multipath structure of propagation described in Section IV, the challenge is to coordinate the radio access, allocate the available radio resources, adjust the radio link parameters and predistort transmitted waveform where appropriate to achieve a well-defined ISAC quality of service. This applies to both radar target localization performance and communication performance, although the two performance measures will not be independent of each other. Although we aim to reuse the available resources (e.g. use the same waveform for communication and sensing) in order to maximize resources efficiency, the goals for performance optimization may be competitive.
@@
The ISAC sensing Key Performance Indicatorss (KPIs) are not yet well defined. On the one hand, as with any radar, the fundamental KPIs are related to localization accuracy and resolution in terms of the number of dynamic targets that ca be resolved and tracked in a certain area. More specifically, this includes also resolvable parameter dimensions like range, Doppler and direction, target acquisition time, coverage and range. Many of the KPI parameters are statistical in nature, such as detection rate, probability of false alarm or missing targets, probability of true tracks and false tracks. Others refer to the detection of target types and shapes. Still others refer to image quality or the ability to detect, analyze and evaluate complex scenarios. Due to the largely ubiquitous availability of wireless sensors and the wide variety of assignable radio resources on the radio and network level, ISAC has the potential to be a highly customizable radar sensor system that can meet various performance requirements. A comprehensive discussion of ISAC KPIs is beyond the scope of this paper. For parameter variance and resolution assessment, the classical parameter is the Cramer Rao lower bound which we apply also for waveform design. Some recent works [ref19, ref50] suggest mutual information (MI) as a common performance metric for the joint radar and communications system. However, there is still a lack of effort in defining new metrics for MIMO radar and MU-MIMO communications co-design problems. Joint optimization, which takes into account the balance between communication and sensing quality of service within the constraints of limited availability and economy of resources, appears to be uncharted territory—but is of great importance for resource scheduling at the network level.
@@ V-A Multistatic Dynamic Target State Vector Estimation
Figure 7: 3D eToF and eDoppler ellipsoids and 3D target state vector. The target is characterized by its instantaneous position vector (here in 2D) and by its differential speed vector with its estimated components. The ISAC sensor nodes are defined by the same two components. But for simplicity only the speed vectors are depicted here. The positions are indicated just by dots.
@@
ISAC aims to estimate the dynamic target state vector, which describes the position, orientation (if possible) and movement of the target in three dimensions. Therefore, we usually need several measurements taken by spatially distributed ISAC radio nodes to determine targets coordinates and movements. These nodes can closely cooperate at the intra-cluster level or more loosely at the inter-cluster level. While in the first case a local, sensor centric coordinate system is mostly applied, we need in the second case a sensor independent global coordinate system since target coordinates are exchanged between different sensor clusters. The sensor centric coordinate system can be Cartesian or polar. While for single station monostatic radars, where range and bearing are measured, the polar coordinate system relative to the radar‘s position is the natural choice, the situation with MS ISAC is somewhat more complicated. Due to the bistatic nature of radio access, the iso-eToF surfaces spanned by the Tx/Rx pairs forming the cluster of closely cooperating ISAC nodes are equipotential ellipsoids, as shown in Figure 7. The same figure also depicts the chosen global cartesian coordinate system of the dynamic state vector, whose static components result from the crossing point of a measured eToF ellipsoids from several Tx / Rx pairs. The dynamic state vector part of moving targets is characterized be their spatial velocity vector and (if available) by higher spatial derivatives, such as acceleration. The components of the bistatic Excess Velocity vector (eVV) target velocity vector (relative to the velocity the moving pair of sensors) is identified from the measured excess Doppler shift and its projection to the respective dynamic sensor state vectors of the sensor nodes. With the static part we may want extended targets to be localized in relation to some single focal point. However, the targets reveal by the first approaching reflection from the target surface relative to the sensor nodes. Multiple bistatic measurements can help to localize the target with respect to its center of gravity by taking advantage of the target related diversity and proper averaging. Moreover, the orientation and shape of the target as well as its inherent movement (micro Doppler) may be of interest, but this would go beyond the scope of this article.
@@
From a set of $L$ independent eToF measurements with transmitter and receiver positions $\bm{r}_{\rm tx_{i}}$, $\bm{r}_{\rm tx_{i}}$, and corresponding eToF observations $\tau_{i}$ and speed of light $c$, we can infer that the position $\bm{r}_{\rm target}$ of a point target has to satisfy
$\left\|\bm{r}_{\rm tx_{i}}-\bm{r}_{\rm target}\right\|_{2}+\left\|\bm{r}_{\rm rx_{i}}-\bm{r}_{\rm target}\right\|_{2}=\tau_{i}c,\quad 1\leqslant i\leqslant L.$ (4)
@@
which corresponds to a system of quadratic equations, which always has a unique solution as soon as $L\geqslant 3$ and if all three ellipses correspond to the same target (c.f. discussion about extended targets above). Similarly, we can derive a quadratic set of equations between the observed excess Doppler (eDoppler) $\alpha_{i}$ at pair of moving ISAC nodes that observe a moving target with velocity $\bm{v}_{\rm target}$. Here, we assume that both target as well as Tx and Rx move with locally constant velocity, which allows us to formulate the relationship between the eDoppler $\alpha_{i}$ and the desired velocity $\bm{v}_{\rm target}$ in Euclidean coordinates via
@@
Here, $\left\langle\cdot,\cdot\right\rangle$} denotes the Euclidean standard 3D inner product. However, this set of equations is only valid, if the eToF $\tau_{i}$ and eDoppler $\alpha_{i}$ can be estimated without error, which in the presence of noise is impossible. Hence, regularization or least-squares approaches need to be considered.
@@
A visualization of the relationship between target position, target velocity and the sensing nodes‘ positions and velocities is given in Figure 7. For each link $i$ we depict the observed ToF $\tau_{i}$ as ellipses, whose foci are the sensing nodes‘ positions and the target is a point on the respective ellipse. Obviously, all the ellipses intersect at the target’s location. Further, at each ISAC link, the target‘s current velocity imposes an observed Doppler shift $\alpha_{i}$ that can be geometrically derived by orthogonally projecting the velocity $\bm{v}_{\rm target}$ onto the line that is perpendicular to the ellipse at the target‘s position. Then after scaling by $\cos{\beta_{i}/2}$, where $\beta_{i}$ denotes the bistatic angle between Tx, target and Rx, we obtain $\alpha_{i}$. Note that for correct calculations all positions must be given in the same global coordinate system. Further, the projected vectors $\alpha_{i}$ as depicted in Figure 7 are drawn to scale for static ISAC nodes. In the case of relative motion, the full extent of (V-A), has to be considered to calculate the $\alpha_{i}$.
@@
While a detailed consideration of the estimation of the target parameters from (4) is beyond the scope of this paper, a general discussion seems worthwhile. Since the equations are nonlinear, the global solution can be ambiguous. This is especially true if single measurements are erroneous. Using surplus measurements for averaging maybe desired. Then a linearization may be helpful which leads to a local linear Taylor series approximation of (4). This is most useful for differential estimation relative to former verified positions. This paves the road to achieve a least squares position estimate by applying the Moore-Penrose pseudo inverse to an overdetermined set of linear equations gained from excess measurements. This procedure takes also advantage from the target related diversity gain as discussed above. However, the choice of bistatic Tx / Rx sensor pairs should be made very careful which makes it an important issue of resource allocation. Not only that the radio access of selected pairs of nodes has to be coordinated and adapted in order to achieve the desired accuracy and resolution of the bistatic measurements. In addition to reducing the variance by averaging, an excess number of measurements can also be used to identify outliers that may occur, for example, due to an obstructed LOS. The sensor fusion of the measurement data must also be carried out. The selection of ISAC nodes also affects data transmission and computation in some fusion centers. The assignment of the nodes depends on their suitability and relevance, which in turn depends on the geometric constellation of the node pairs in relation to the target and other interacting objects. This has influence not only to the SNR of the estimates. There is also an error propagation mechanism known from satellite navigation as Geometric Dilution of Precision (GDoP) as briefly indicated in Figure 8.

### v2

@@ IV-A Multipath Propagation Channel and Multipath Exploitation
Understanding of multipath propagation issues is crucial to the performance of mobile radio and are critical to system design. Multipath is responsible for channel dispersion in the delay, Doppler, and angular domains. It leads to temporal variability, fast and slow fading. It has been considered the most limiting feature of system performance that needs to be mitigated by equalization, precoding and link adaptation in different domains. Despite a certain diversity gain advantage, multipath originally was seen as harmful for high data rates and reliable communication. The situation has dramatically changed with the advent of MIMO radio access, which has shown that multipath can improve communication and enhance capacity in complicated, multipath-rich environments. Also in radio localization, multipath may turn from a foe to a friend if correctly exploited [ref55]. Therefore, measurement and model based propagation studies in rich scattering environments are of outstanding importance also for ISAC system design [ref3]. Although the topic would deserve more attention, we can give only a concise overview on the most relevant propagation effects to support understanding of basic MS-ISAC design issues.
@@
The consideration of the propagation channel for radar localization differs remarkably from the situation in communication. Generally speaking, in radar the desired information is contained in the geometric structure of the propagation, while in communication the information is encoded in the unknown transmitted symbols, which have to be received correctly. With ISAC, where communication and sensing functions are integrated, we even have a mixture of both which casts different views on the influence of propagation. Figure 6 depicts the geometric structure of a single bistatic multipath propagation channel. In communication, the total energy per symbol sent and received is used for symbol recognition, hence for maximizing mutual information. In ISAC only those paths that are routed via the target can contribute to target detection and localization. However, the LOS path between Tx and Rx (which is often misleadingly referred to simply as “interference”) is also important. In addition to its contribution to communication, it helps for Tx2Rx frame synchronization, submits a reference for differential eToF and eDoppler estimation in bistatic setups and correlation reference signal recovery, which is necessary for CPCL target estimation as described section III. In contrast, all other propagation paths that are not routed via the target are useless for radar sensing, perhaps disturbing and therefore referred to as clutter. The role of the target channel, which consists of all multiple paths that are routed via the target, can be thought of as a relay channel as highlighted in Figure 6.
@@
The target is most clearly identified by the two LOS paths (Tx-to-target and target-to-Rx in Figure 6), which we call the Bistatic Target Path (BTP). However, the simple triangular LOS-related model would be invalid if one or more of the three LOS paths are blocked. While the blocked direct Tx2Rx LOS must be given special consideration since it delivers the correlation reference, there are several possibilities to exploit target related multipath. In general, the associated multipath links that are also routed via the target additionally contribute to targets illumination and visibility and may reveal relevant geometric information for localization even if the respective LOS is missing. Spatial precoding and dedicated estimation measures would be necessary to make use of this multipath interaction (“multipath exploitation”), may it be enhanced visibility through spatial multipath focusing (increased illumination power) on the target or enabling target visibility even when the target LOS to Rx is obstructed. Here we can only give a very concise overview highlighting effects and indicating respective measures to mitigate or exploit multipath.
@@
Proactive exploitation of multipath goes even further. While precoding and associated resource allocation measures are shortly referred to in the next section, a detailed discussion about multipath exploiting procedures is beyond the scope of this paper. Roughly speaking, there seem to be two classes of multipath exploiting estimation methods. One is called “finger printing”. It is based on calculating some position related features from observed ISAC data. In the most simple case, this can be received power which undergoes spatial selective fading. However, differential delays, Doppler or directions that can be better attributed to relative target position and movement can be also used. While explicit geometrical calculation of position from these parameters may not be possible in multipath rich environment, machine learning can be applied to identify targets position from measured fingerprints. In addition, reference information calculated from a ray tracing simulation of the scenario (a “digital twin”) can serve as a learning data base [ref56, ref57].
Another approach is to explicitly embeddi a multipath propagation data model based on ray tracing into the position estimation loop. Assuming we know the positions of the dominant interaction points of the environment, the associated reflections can be used as further (or auxiliary) spatially orthogonal measurements like additional illuminations generated by virtual anchor nodes. This inverse model-based estimator effectively traces back the observed specular data to the unknown target position [ref51]. The advantage relative to finger printing is less learning effort which is required (the explicit knowledge about the structure of the environment is directly applied) and we have more control about which interacting objects are included. This may be also important if “artificial interactions” generated by Reconfigurable Intelligent Surfaces (RISs) or by amplify-and-forward relays are used.
Yet another approach that enables target localization with obstructed LOS from Tx to or from the target to Rx (e.g. if the target gets lost when tracking) makes use of knowledge about the effect of obstruction and especially if we explicitly know the position of the obstructing object [ref58, ref59]. Even when the target itself is obscuring the Tx2Rx LOS we can detected the presence of the target when it moves through the LOS between Tx and Rx (forward scattering radar [ref60]).
@@
where $H_{\rm los}$, $H_{\rm dti}$, $H_{\rm dtr}$, $H_{\rm sti}$, $H_{\rm str}$, and $H_{\rm clt}$ denote the FRFs for the LOS, direct target illumination, direct target response, scattered target illumination, scattered target response, and clutter, respectively. Clutter summarizes all contributions that are not related to the target (except LOS). Whether the scattered target illumination and response should be considered a clutter depends on whether it is exploited or not, i.e., if it constitutes unwanted interference.
@@ IV-B Multistatic Target Reflectivity and Diversity Gain
The target’s contribution to the end-to-end multipath routing from transmitter to receiver is described by its Multistatic Reflectivity Function (MRF), which is the bistatic transfer function of all pairs of incoming and outgoing BTP waves including target-related LOS and multipath. Therefore, the target deserves special attention in ISAC propagation measurement and modeling. A complete MRF description would be based on 2x2 full polarimetric bistatic angular scanning. The MRF measurement range described in [ref3] uses wideband illumination in order to characterize a target’s inherent time variability (e.g. because of rotating wheels or propellers in case of cars or drones or moving arms and legs in case of walking persons), which will cause characteristic micro Doppler patterns [ref70]. Together with the bistatic spread in delay and angle, we obtain a full MRF description that enables advanced propagation studies, recognition of extended targets and high performance ray optical modeling in multipath rich environments. Note that these characteristics are distance dependent as they are influenced by spherical wavefronts.
@@
At greater distances and lower bandwidths (which does not allow to resolve the shape of the target), the advantage of distributed MIMO radar shifts to spatial diversity, which affects the probability of target detection. The use of spatial diversity for mitigation of fading is already well known from MIMO communication where it is related to the signal at the receiver antenna position. In distributed MIMO radar it is related the target. Since radar targets are usually larger than several wavelengths of the radio carrier (“electrically large targets”) and complex in terms of many structural details of small size, even cavities, the scattered field is composed of contributions from a great number of seemingly distinct scattering centers, with the resulting superposition being a complex circular symmetric Gaussian variable (iid in real and imaginary part). Therefore, the reflected power of the target (related to its Radar Cross Section (RCS)) follows a $\chi^{2}$ distribution with two degrees of freedom. This applies as long as the multiple reflections make a similar contribution (in terms of power) to the resulting target backscattering. In case we have a dominating (specular) contribution, the distribution will change to a non-central chi-square distribution (a squared Rician distribution). Moreover, the reflected power maybe subject to a slow variability over time. These scintillations are responsible for signal fading, which has influence to target detection statistics. In radar technology, these statistical variations are combined in so called Swerling’s target models [ref32]. In mobile radio communications, the same phenomenon is known as small scale and large scale fading.
@@
Diversity gain increases the probability of detection in case of multiple observations [ref10c]. It depends on the decorrelation of the combined observations and can be characterized as cooperative diversity or macro-diversity. Therefore, distributed MIMO radar is often referred to as statistical MIMO radar to distinguish it from MIMO radar with coherent signal combining. The possible diversity gain depends of course on the number of independent observations, which in the best case can be up to $N^{2}$ (see section III-A). A welcome side effect is that the claimed distributed spatial structure is also advantageous for estimating the 3D dynamic state vector which is described in section V. The influence of the aspect angle is also evident when we consider special target shapes, such as those targets that attempt to conceal themselves by minimizing their monostatic RCS, also known as stealth targets. At the same time, we have to emphasize a very specific spatial constellation of Tx and Rx, the so-called forward scattering case where Tx and Rx face each other from opposite directions and the target is in between. Surprisingly enough, this situation has many advantages for target detection as the received power may undergo a clear increase because of the creeping wave along the targets surface and the resulting lens effect. For higher frequencies (mmWave) and depending on the distance between target and antennas, also shadowing may occur. The latter case may become relevant for short distance sensing in car-to-car scenarios.
@@
While the combined $\chi^{2}$ distributions apply to multiple observations of electrically large targets, we need to distinguish the case of extended targets, which are resolved in the delay domain and are therefore represented by a small number of neighboring range bins. For a given target, this situation depends on the bandwidth of the ISAC waveform. The more bandwidth we have, the more range bins we can resolve – for the same target. Since the observed reflectivity response of the neighboring range bins is not independent because it originates from the same target seen from the same direction, we can try to maximize received power, respectively the detection probability, by matching the transmit signal to the target response (target matched filter).
@@ V MS ISAC Access and Estimation
In Section III, we describe a very general architecture of MS MIMO ISAC systems. According to the multipath structure of propagation described in Section IV, the challenge is to coordinate the radio access, allocate the available radio resources, adjust the radio link parameters, and predistort the transmitted waveform where appropriate to achieve a well-defined ISAC quality of service. This applies to both radar target localization performance and communication performance, although the two performance measures will not be independent of each other. Although we aim to reuse the available resources (e.g., use the same waveform for communication and sensing) in order to maximize resource efficiency, the goals for performance optimization may be competitive.
@@
The ISAC sensing Key Performance Indicators (KPIs) are not yet well defined. On the one hand, as with any radar, the fundamental KPIs are related to localization accuracy and resolution in terms of the number of dynamic targets that can be resolved and tracked in a certain area. More specifically, this also includes resolvable parameter dimensions such as range, Doppler and direction, target acquisition time, and coverage. Many of the KPI parameters are statistical in nature, such as detection rate, probability of false alarm or missing targets, probability of true tracks and false tracks. Others refer to the detection of target types and shapes. Still others refer to image quality or the ability to detect, analyze, and evaluate complex scenarios. Due to the largely ubiquitous availability of wireless sensors and the wide variety of assignable radio resources on the radio and network level, ISAC has the potential to be a highly customizable radar sensor system that can meet various performance requirements. A comprehensive discussion of ISAC KPIs is beyond the scope of this paper. For parameter variance and resolution assessment, the classical parameter is the Cramér-Rao Bound (CRB), which we also apply for waveform design. Some recent works [ref19, ref50] suggest Mutual Information (MI) as a common performance metric for the joint radar and communications system. However, there is still a lack of effort in defining new metrics for MIMO radar and MU MIMO communications co-design problems. Joint optimization, which takes into account the balance between communication and sensing quality of service within the constraints of limited availability and economy of resources, appears to be uncharted territory—but is of great importance for resource scheduling at the network level.
@@ V-A Multistatic Dynamic Target State Vector Estimation
Figure 7: 3D eToF and eDoppler ellipsoids and 3D target state vector. The target is characterized by its instantaneous position vector (here in 2D) and by its differential speed vector with its estimated components. The ISAC sensor nodes are defined by the same two components. But for simplicity, only the speed vectors are depicted here. The positions are indicated just by dots.
@@
ISAC aims to estimate the dynamic target state vector, which describes the position, orientation (if possible), and movement of the target in three dimensions. Therefore, we usually need several measurements taken by spatially distributed ISAC radio nodes to determine targets’ coordinates and movements. These nodes can closely cooperate at the intra-cluster level or more loosely at the inter-cluster level. While in the first case a local, sensor-centric coordinate system is mostly applied, we need in the second case a sensor-independent global coordinate system since target coordinates are exchanged between different sensor clusters. The sensor-centric coordinate system can be Cartesian or polar. While for single station monostatic radars, where range and bearing are measured, the polar coordinate system relative to the radar‘s position is the natural choice, the situation with MS-ISAC is somewhat more complicated. Due to the bistatic nature of radio access, the iso-eToF surfaces spanned by the Tx / Rx pairs forming the cluster of closely cooperating ISAC nodes are equipotential ellipsoids, as shown in Figure 7. The same figure also depicts the chosen global Cartesian coordinate system of the dynamic state vector, whose static components result from the crossing point of a measured eToF ellipsoids from several Tx / Rx pairs. The dynamic state vector part of moving targets is characterized by their spatial velocity vector and (if available) by higher spatial derivatives, such as acceleration. The components of the bistatic Excess Velocity vector (eVV) target velocity vector (relative to the velocity of the moving pair of sensors) are identified from the measured excess Doppler shift and its projection to the respective dynamic sensor state vectors of the sensor nodes. With the static part, we may want extended targets to be localized in relation to some single focal point. However, the targets are revealed by the first approaching reflection from the target surface relative to the sensor nodes. Multiple bistatic measurements can help to localize the target with respect to its center of gravity by taking advantage of the target-related diversity and proper averaging. Moreover, the orientation and shape of the target, as well as its inherent movement (micro Doppler) may be of interest, but this would go beyond the scope of this article.
@@
From a set of $L$ independent eToF measurements with transmitter and receiver positions $\bm{r}_{\rm tx_{i}}$, $\bm{r}_{\rm rx_{i}}$, and corresponding eToF observations $\tau_{i}$ and speed of light $c$, we can infer that the position $\bm{r}_{\rm target}$ of a point target has to satisfy
$\left\|\bm{r}_{\rm tx_{i}}-\bm{r}_{\rm target}\right\|_{2}+\left\|\bm{r}_{\rm rx_{i}}-\bm{r}_{\rm target}\right\|_{2}=\tau_{i}c,\quad 1\leqslant i\leqslant L$ (4)
@@
which corresponds to a system of quadratic equations, which always has a unique solution as soon as $L\geqslant 3$ and if all three ellipses correspond to the same target (c.f. discussion about extended targets above). Similarly, we can derive a quadratic set of equations between the observed excess Doppler (eDoppler) $\alpha_{i}$ at a pair of moving ISAC nodes that observe a moving target with velocity $\bm{v}_{\rm target}$. Here, we assume that both target as well as Tx and Rx move with locally constant velocity, which allows us to formulate the relationship between the eDoppler $\alpha_{i}$ and the desired velocity $\bm{v}_{\rm target}$ in Euclidean coordinates via
@@
Here, $\left\langle\cdot,\cdot\right\rangle$} denotes the Euclidean standard 3D inner product. However, this set of equations is only valid if the eToF $\tau_{i}$ and eDoppler $\alpha_{i}$ can be estimated without error, which in the presence of noise is impossible. Hence, regularization or least-squares approaches need to be considered.
@@
A visualization of the relationship between target position, target velocity, and the sensing nodes‘ positions and velocities is given in Figure 7. For each link $i$ we depict the observed ToF $\tau_{i}$ as ellipses, whose foci are the sensing nodes‘ positions, and the target is a point on the respective ellipse. Obviously, all the ellipses intersect at the target’s location. Further, at each ISAC link, the target‘s current velocity imposes an observed Doppler shift $\alpha_{i}$ that can be geometrically derived by orthogonally projecting the velocity $\bm{v}_{\rm target}$ onto the line that is perpendicular to the ellipse at the target‘s position. Then after scaling by $\cos{\beta_{i}/2}$, where $\beta_{i}$ denotes the bistatic angle between Tx, target and Rx, we obtain $\alpha_{i}$. Note that for correct calculations, all positions must be given in the same global coordinate system. Further, the projected vectors $\alpha_{i}$ as depicted in Figure 7 are drawn to scale for static ISAC nodes. In the case of relative motion, the full extent of (V-A) has to be considered to calculate the $\alpha_{i}$.
@@
While a detailed consideration of the estimation of the target parameters from (4) is beyond the scope of this paper, a general discussion seems worthwhile. Since the equations are nonlinear, the global solution can be ambiguous. This is especially true if single measurements are erroneous. Using surplus measurements for averaging may be desired. Then a linearization may be helpful, which leads to a local linear Taylor series approximation of (4). This is most useful for differential estimation relative to former verified positions. This paves the road to achieve a least squares position estimate by applying the Moore-Penrose pseudo inverse to an overdetermined set of linear equations gained from excess measurements. This procedure also takes advantage of the target-related diversity gain as discussed above. However, the choice of bistatic Tx / Rx sensor pairs should be made very carefully, which makes it an important issue of resource allocation. Not only does the radio access of selected pairs of nodes have to be coordinated and adapted in order to achieve the desired accuracy and resolution of the bistatic measurements. In addition to reducing the variance by averaging, an excess number of measurements can also be used to identify outliers that may occur, for example, due to an obstructed LOS. The sensor fusion of the measurement data must also be carried out. The selection of ISAC nodes also affects data transmission and computation in some fusion centers. The assignment of the nodes depends on their suitability and relevance, which in turn depends on the geometric constellation of the node pairs in relation to the target and other interacting objects. This has an influence not only on the SNR of the estimates. There is also an error propagation mechanism known from satellite navigation as Geometric Dilution of Precision (GDoP) as briefly indicated in Figure 8.